The uncertainty in the momentum Δp of a golf ball stroked by a player which is traveling at 50m/s is  1×10−6 of its momentum. What is its uncertainty in position Δx? Mass of the golf ball was measured as 250 g 



1
Expert's answer
2021-03-24T19:41:13-0400

According to the Heisenberg's uncertainty principle


"\Delta p\Delta x \ge\dfrac{\hbar}{2}"

where "\hbar \approx 1.05\times 10^{-34}J\cdot s" is the reduced Planck constant. According to the task


"\Delta p = 10^{-6}\cdot mv"

where "m = 250g = 0.25kg" is the mass of the ball, "v = 50m\/s" is its speed.

Thus, the smalles possible uncertainty in position is:


"\Delta x = \dfrac{\hbar}{2\Delta p} = \dfrac{\hbar}{2\cdot 10^{-6}\cdot mv}\\
\Delta x = \dfrac{ 1.05\times 10^{-34}}{2\cdot 10^{-6}\cdot 0.25\cdot 50} \approx 4.2\times 10^{-30}m"

Answer. "4.2\times 10^{-30}m",